OpenStudy (anonymous):

How to factor d^2+10d+9

OpenStudy (anonymous):

you have to find the solutions to d^2 +10d + 9 = 0

OpenStudy (anonymous):

d^2+9d+d+9 d(d+9)+1(d+9) (d+9)(d+1) i guess its enough there is no need to put it equal to 0 as u r asking to factorize not to find the roots

OpenStudy (anonymous):

your solution is more simple, but i'm afraid it must me hard sometimes (but i gess it's enough for tayler).

OpenStudy (anonymous):

you could use the formula: d1 = [-10+sqrt(10^2 - 4*9)]/2 d1 = (-10+sqrt64)/2 d1 = (-10+8)/2 d1 = -1 d2 = (-10-8)/2 d2 = -9

OpenStudy (anonymous):

Then, knowing the solutions of the equation, d1 and d2, you can factorize: (d - d1)(d - d2) = (d + 1)(d + 9)

OpenStudy (anonymous):

hey niwi are u on facebook

OpenStudy (anonymous):

no